# Autoencoding-Free Context Compression for LLMs via Contextual Semantic Anchors

### v3

@@ Abstract
Context compression presents a promising approach for accelerating large language model (LLM) inference by compressing long contexts into compact representations. Current context compression methods predominantly rely on autoencoding tasks to train context-agnostic compression tokens to compress contextual semantics. While autoencoding tasks enable compression tokens to acquire compression capabilities, compression via autoencoding tasks creates a fundamental mismatch: the models are optimized for reconstruction that diverge from actual downstream tasks, thereby weakening the features more beneficial for real-world usage. We propose Semantic-Anchor Compression (SAC), a novel method that shifts from autoencoding task based compression to an architecture that is equipped with this compression capability a priori. Instead of training models to compress contexts through autoencoding tasks, SAC directly selects so-called anchor tokens from the original context and aggregates contextual information into their key-value (KV) representations. By deriving representations directly from the contextual tokens, SAC eliminates the need for autoencoding training. To ensure compression performance while directly leveraging anchor tokens, SAC incorporates two key designs: (1) anchor embeddings that enable the compressor to identify critical tokens, and (2) bidirectional attention modification that allows anchor tokens to capture information from the entire context. Experimental results demonstrate that SAC consistently outperforms existing context compression methods across various compression ratios. On out-of-distribution evaluation using MRQA, SAC achieves 1 EM improvement at 5x compression over strong baselines, with increasing advantages at higher compression ratios.
++footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
The expanding scope of large language models (LLMs) to tasks like processing long documents (Liu et al., 2024b; Li et al., 2024; Duan et al., 2025), maintaining multi-turn dialogue coherence (Zhang et al., 2025; Yi et al., 2025; Guan et al., 2025), and generating responses grounded in extensive external knowledge (Lewis et al., 2020; Karpukhin et al., 2020; Huang et al., 2025) necessitates the incorporation of vast contexts into the model input. However, directly processing such extremely long contexts is fraught with challenges, including prohibitive computational costs, significant inference latency, and performance degradation, largely caused by the “lost-in-the-middle” phenomenon (Liu et al., 2024a).
To address these challenges, recent studies have proposed context compression (Chang et al., 2024; Li et al., 2025a), a technique that typically appends special tokens (i.e. compression tokens) to the end of the context and leverages the LLM’s causal attention mechanism to compress contextual information into a compact representation within these tokens. Once this compact representation is obtained, the LLM can generate responses conditioned on it, rather than being conditioned on the entire original context. This significant reduction in context length leads to substantial decreases in both inference time and GPU memory consumption. While effective, these approaches face a key limitation: the compression tokens are randomly initialized and lack inherent semantic information. To compensate, they typically rely on extensive pretraining on both autoencoding (AE) and language modeling (LM) tasks (illustrated in Figure 1) to endow the compression tokens with the ability to carry contextual information. While AE task has shown to be necessary in ICAE (Ge et al., 2024) since compression tokens lack context-relevant semantics, the AE task requires the compressed representation to reconstruct all tokens in the context, even low-information tokens. This reliance on a suboptimal and costly pretraining stage raises a critical research question: Is it possible to design a compression architecture that inherently understands context without a demanding AE phase?
@@
To answer this question, this work introduces Semantic-Anchor Compression (SAC) (Figure 2), a novel architecture for the context compression task. Instead of appending new special tokens and requiring extensive autoencoding pretraining to learn their representations, SAC directly selects representative tokens from the original context to act as ‘anchor tokens’ for compression. By leveraging these semantically meaningful anchors from the input itself, SAC incorporates natural semantic priors that obviate the need for autoencoding pretraining. To signify their special role, these selected tokens are then augmented with dedicated ‘anchor embeddings’, enabling the LLM to distinguish them from regular tokens. Furthermore, to enhance their compression capabilities, we modify the standard causal attention to a bidirectional attention mechanism. This allows anchor tokens to access information from the entire context, rather than being restricted to only preceding tokens. These modifications collectively foster a more effective context compression by providing anchor tokens with both distinct representations and comprehensive contextual awareness. Empirically, we test SAC on the MRQA (Fisch et al., 2019a) dataset and confirm that it outperforms existing strong context compression baselines. For example, compared to 500xCompressor (Li et al., 2025b) at 5x compression, the average exact match (EM) improves from 25.4 to 32.3. Results show that 1) our proposed method improves more in absolute accuracy over strong baselines on more challenging high compress ratio scenarios 2) our proposed architecture achieves its best performance in a simpler training setting without autoencoding training arguably because the anchor tokens already contain enough information about the original context. Our analysis reveals that SAC’s compressed representations more closely resemble original context token KVs in feature space, so that LLMs performing inferenece can arguably better understand them.
Figure 2: The difference between SAC and previous methods. While previous methods (a) compress contextual information into context-agnostic special tokens (referred to as compression tokens), SAC (b) compresses the context directly into the original contextual tokens themselves. Here, $\tilde{M}$ can represent either the output from the final layer of the LLM or the Key-Value pairs, being later on used as compressed representations for LLM inference.
@@ 2.1 Compression method
Many methods focus on reducing prompt lengths. CC (Wingate et al., 2022) utilizes contrastive learning to compress specific natural language prompts into shorter and unique soft prompt tokens. However, it cannot generalize to unseen prompts and requires retraining for new prompts. GIST (Mu et al., 2023) compresses original prompts into KV values through finetuning and can handle arbitrary unseen contexts. AutoCompressor (Chevalier et al., 2023) recursively combines compressed vectors with sub-prompts and aggregates all compressed vectors to construct the final representation, enabling compression of longer contexts. However, both GIST and AutoCompressor require finetuning the LLMs performing inference (referred to later as target LLM), which may affect LLMs’ original capabilities.
ICAE (Ge et al., 2024) formulates context compression as training a general encoder that compresses contexts into compact representations understandable by target LLMs without finetuning. To achieve this, ICAE introduces autoencoding tasks and performs joint pretraining with language modeling tasks, followed by finetuning on downstream tasks. 500xCompressor (Li et al., 2025b) improves upon ICAE by replacing the compression carrier from the last layer output of compression tokens with KV values at each layer, achieving higher compression ratios. EPL (Zhao et al., 2025) identifies that ICAE and 500xCompressor neglect the impact of positional encoding and proposes distributing compression token position IDs uniformly across the entire context rather than placing them at the end. However, these methods still rely on autoencoding tasks to endow the compression tokens with the ability to carry contextual information.
Another category of prompt compression methods is based on token selection, which selects representative tokens from contexts based on token importance. SelectiveContext (Li et al., 2023), LLMLingua (Jiang et al., 2023), and LongLLMLingua (Jiang et al., 2024) employ causal small language models to evaluate token importance based on information entropy. LLMLingua-2 (Pan et al., 2024a) distills a token classifier to compute the probability of each token to be preserved. These works demonstrate that LLMs can understand original contexts using a small number of representative tokens. However, they do not perform compressed tokens training which limits the usability of the selected tokens by target LLMs. Our proposed SAC can be seen as a combination of token selection methods and compressed token training methods: it derives and train compressed representations that are based on tokens selected directly from the context and indeed is compatible with the token selection methods above.
@@ 2.2 Bidirectional Attention
Recent studies have shown that, removing the decoder’s unidirectional causal constraint and introducing bidirectional attention can effectively enhance the model’s representational capacity (Wang et al., 2020). For instance, NV-Embed (Lee et al., 2025) replaces causal attention with bidirectional attention during contrastive training, achieving strong performance on general text embedding and dense vector retrieval tasks. LLM2Vec (BehnamGhader et al., 2024), by enabling bidirectional attention alongside masked next-token prediction, significantly improves the model’s ability to capture global semantics in text embedding tasks. These works indicate that bidirectional attention is advantageous for acquiring global semantic information and robust contextual representations. However, its effectiveness in text compression tasks remains underexplored. Motivated by these findings, we incorporate bidirectional attention into the compressor to enhance contextual modeling during the compression phase.
@@ 3.1 Task Formulation
Additionally, during finetuning, the question answering loss $\mathcal{L}_{\mathbf{QA}}$ enhances the ability of $\tilde{M}$ to extract information that is potentially relevant for downstream tasks (e.g. QA). Since the encoder operates without knowledge of what questions might be asked later, it learns to identify and preserve information that is likely to be queried, enabling accurate answer generation $A=(a_{1},a_{2},\ldots,a_{|A|})$ when presented with subsequent questions $Q=(q_{1},q_{2},\ldots,q_{|Q|})$, as shown in Figure 1c; mathematically, $\mathcal{L}_{\mathbf{QA}}=-\log P(A|\tilde{M},Q)$.
@@ 3.2 Semantic-Anchor Compressor
A key distinction between our approach SAC and previous methods is that we derive compressed representations directly from selected context tokens, as shown in Figure 2. This involves selecting a subset of tokens from context $C$ as anchor tokens $S\subseteq C$. We believe that a good selection strategy benefits SAC. Following EPL, our default strategy divides the entire context $C$ into $|S|$ chunks and selects the middle token from each chunk. This setting helps maximize coverage of context $C$. As illustrated in Figure 3a, selected tokens $c_{i}\in S$ are enhanced with anchor embeddings $e_{A}$, yielding an embedding sequence $E=(e_{1},e_{2},\ldots,e_{|C|})$:
@@
where $\mathbf{1}_{c_{i}\in S}$ is an indicator function that equals 1 when $c_{i}\in S$ and 0 otherwise. Following previous works, we employ a LLM with LoRA parameters $\theta_{LoRA}$ as the compressor: $\tilde{M}=\mathcal{E}(C)=\mathbf{LLM}(E|\theta_{LoRA})$.
While using original tokens from the context avoids learning compressed tokens from scratch and potentially improves the learning efficiency. We notice that because the encoder uses causal attentions, the anchor tokens $S$ do not have visibility to the full sentence, limiting its representation power. Hence we modify the LLM from using causal attention to use bidirectional attention (see Figure 3b), enhancing the LLM’s encoding capability. $\tilde{M}$ can be either the output of anchor tokens from the LLM’s final layer or the Key-Value pairs from each layer. Following 500xCompressor, we use Key-Value pairs as the compressed representation $\tilde{M}$. During pretraining, we only use $\mathcal{L}_{\mathbf{LM}}$ and do not use $\mathcal{L}_{\mathbf{AE}}$ to train the compressor. Following previous work, we use $\mathcal{L}_{\mathbf{QA}}$ for finetuning.
@@ 4.1 EXPERIMENTAL SETTING
Dataset. For continued pretraining, we utilize the large-scale corpus  SlimPajama-6B (Soboleva et al., 2023). During finetuning and evaluation, we employ the standard MRQA (Fisch et al., 2019b) question-answering dataset, which consolidates multiple QA tasks and standardizes them into a unified format. We evaluate SAC on both test sets, namely in-domain (ID) and out-of-domain (OOD), to comprehensively assess its in-distribution fitting ability and cross-domain generalization performance.
@@
Implementation Details. SAC utilizes Llama-3.2-1B (Grattafiori & Dubey, 2024) as both the encoder and target LLM. The encoder is equipped with trainable LoRA (Hu et al., 2022) adapters (rank = 128, $\alpha=256$), while the target LLM parameters remain frozen. For each context, we partition it into sub-contexts of 510 tokens each. The compressor compresses each sub-context into a sub-compressed representation, and subsequently concatenates these sub-compressed representations to form the complete compressed representation. The number of anchor tokens $|S|=\lfloor L/r\rfloor$ is determined by the compression ratio $r$ and the length of the sub-context $L$. We train all models in two stages: pretraining for 20,000 optimization steps followed by finetuning for an additional 20,000 steps, both conducted with a batch size of 16. Complete hyperparameter configurations are provided in Appendix A.

### v4

@@ Abstract
Context compression is an advanced technique that accelerates large language model (LLM) inference by converting long inputs into compact representations. Existing methods primarily rely on autoencoding tasks to train special compression tokens to represent contextual semantics. While autoencoding tasks enable compression tokens to acquire compression capabilities, we remark that such capabilities potentially conflict with actual downstream task requirements, prevent the models from learning the features more beneficial for real-world usage. Based on this observation, we propose Semantic-Anchor Compression (SAC), a novel method that shifts from autoencoding task based compression to an architecture that is equipped with this compression capability a priori. Instead of training models to compress contexts through autoencoding tasks, SAC directly selects so-called anchor tokens from the original context and aggregates contextual information into their key-value (KV) representations. To ensure that anchors can effectively collect information, SAC introduces two key designs: (1) anchor embedding, a learnable embedding vector attached to the selected anchor tokens to mark compression carriers and (2) bidirectional attention modification, which enables anchor tokens to integrate information from the entire context. Experimental results show that SAC consistently outperforms existing context compression methods across different compression ratios and model sizes on question-answering and long-context summarization tasks. Our data, model and code have been released at https://github.com/lx-Meteors/SAC.
@@ 1 Introduction
The expanding scope of large language models (LLMs) to tasks such as processing long documents (Liu et al., 2024b; Li et al., 2024; Duan et al., 2025; Zhao et al., 2026), maintaining multi-turn dialogue coherence (Zhang et al., 2025; Yi et al., 2025; Guan et al., 2025), and generating responses grounded in extensive external knowledge (Lewis et al., 2020; Karpukhin et al., 2020; Huang et al., 2025) necessitates the incorporation of vast contexts into the model input. However, directly processing such extremely long contexts is fraught with challenges, including prohibitive computational costs, significant inference latency, and performance degradation, largely caused by the “lost-in-the-middle” phenomenon (Liu et al., 2024a).
To address these challenges, recent studies have proposed context compression (Chang et al., 2024; Li et al., 2025a; Lv et al., 2026; Tang et al., 2026a; b), a technique that typically appends special tokens (i.e. compression tokens) to the end of the context and leverages the LLM’s causal attention mechanism to compress contextual information into a compact representation within these tokens. Once this compact representation is obtained, the LLM can generate responses conditioned on it, rather than being conditioned on the entire original context. The reduction in context length leads to substantial decreases in both inference time and GPU memory consumption. While effective, these approaches face a key limitation: the compression tokens are randomly initialized and lack inherent semantic information. To address this limitation, prior works (Ge et al., 2024; Wang et al., 2024; Li et al., 2025b; Zhao et al., 2025; Tang et al., 2025b) typically rely on extensive pretraining on autoencoding (AE) and language modeling (LM) tasks (as shown in Fig. 1) to endow compression tokens with the ability to carry contextual information. However, while the model learns to reconstruct contexts from compression tokens through the AE task, such reconstruction objective can misalign with actual objectives (e.g. QA from compressd tokens) that downstream requires (we analyse such misalignment in subsection 5.1). Thus, this reliance on a suboptimal and costly pretraining stage raises a critical research question: is it possible to design a compression architecture that inherently understands context without a demanding AE phase?
@@
To answer this question, our work introduces Semantic-Anchor Compression (SAC) (Figure 2), a novel architecture for the context compression task. Instead of appending new special tokens and requiring extensive autoencoding pretraining to learn their representations, SAC directly selects representative tokens from the original context to act as “anchor tokens” for compression. By leveraging these semantically meaningful anchors from the input itself, SAC incorporates natural semantic priors that obviate the need for autoencoding pretraining. To mark their special role, these selected tokens are augmented with dedicated “anchor embeddings”, enabling the LLM to distinguish them from regular tokens. Furthermore, to enhance their compression capabilities, we modify the standard causal attention to a bidirectional attention mechanism. This allows anchor tokens to access information from the entire context, rather than being restricted to only preceding tokens. These modifications collectively foster a more effective context compression by providing anchor tokens with both distinct representations and comprehensive contextual awareness. To evaluate the effectiveness of SAC, we conduct experiments on question answering tasks and long-context summarization tasks, where it consistently outperforms strong baselines. Results also show that 1) our proposed method consistently improves over strong baselines across different compress ratios, 2) our proposed architecture achieves its best performance in a simpler training setting without autoencoding training, arguably because the anchor tokens already contain enough information about the original context. Our analysis reveals that SAC’s compressed representations more closely resemble original context token KVs in feature space, which arguably enables LLMs to better understand them during inference.
Figure 2: The difference between SAC and previous methods. While previous methods (a) compress contextual information into dedicated special tokens (referred to as compression tokens), SAC (b) compresses the context directly into the original contextual tokens themselves. Here, $\tilde{M}$ can represent either the output from the final layer of the LLM or the key-value pairs across all transformer layers, which are later used as compressed representations for LLM inference.
@@ 2.1 Compression method
Many methods focus on reducing prompt lengths to improve LLM inference speed. CC (Wingate et al., 2022) utilizes contrastive learning to compress specific natural language prompts into shorter and unique soft prompt tokens. However, it cannot generalize to unseen prompts and requires retraining for new prompts. GIST (Mu et al., 2023) compresses original prompts into KV values through finetuning and can handle arbitrary unseen contexts. AutoCompressor (Chevalier et al., 2023) recursively combines compressed vectors with sub-prompts and aggregates all compressed vectors to construct the final representation, enabling compression of longer contexts. However, both GIST and AutoCompressor require finetuning the LLMs (referred to later as target LLMs) to perform inference, which may affect LLMs’ original capabilities.
ICAE (Ge et al., 2024) formulates context compression as training a general encoder that compresses contexts into compact representations understandable by target LLMs without finetuning target LLMs. To train the encoder, ICAE introduces autoencoding tasks and performs joint pretraining with language modeling tasks, followed by finetuning on downstream tasks. 500xCompressor (Li et al., 2025b) improves upon ICAE by replacing the compression carrier from the last layer output of compression tokens with KV values at each layer, achieving higher compression ratios. EPL (Zhao et al., 2025) identifies that ICAE and 500xCompressor neglect the impact of positional encoding and proposes distributing compression token position IDs uniformly across the entire context rather than placing them at the end. However, these methods still rely on autoencoding tasks to endow the compression tokens with the ability to carry contextual information.
Another category of prompt compression methods is based on token selection, which selects representative tokens from contexts based on token importance. SelectiveContext (Li et al., 2023), LLMLingua (Jiang et al., 2023), and LongLLMLingua (Jiang et al., 2024) employ causal small language models to evaluate token importance based on information entropy. LLMLingua-2 (Pan et al., 2024a) distills a token classifier to compute the probability of each token to be preserved. PerceptionCompressor (Tang et al., 2025a) preserves critical information at the token level while removing tokens that distract the LLM’s attention. These works demonstrate that LLMs can understand original contexts using a small number of representative tokens. However, they do not perform compressed tokens training which limits the usability of the selected tokens by target LLMs. Our proposed SAC can be seen as a combination of token selection methods and compressed token training methods: it derives and train compressed representations that are based on tokens selected directly from the context and indeed is compatible with the token selection methods above.
@@ 2.2 Bidirectional Attention
Recent studies have shown that, removing the decoder’s unidirectional causal constraint and introducing bidirectional attention can effectively enhance the model’s representational capacity (Wang et al., 2020). For instance, NV-Embed (Lee et al., 2025) replaces causal attention with bidirectional attention during contrastive training, achieving strong performance on general text embedding and dense vector retrieval tasks. LLM2Vec (BehnamGhader et al., 2024), by enabling bidirectional attention alongside masked next-token prediction, significantly improves the model’s ability to capture global semantics in text embedding tasks. These works indicate that bidirectional attention is advantageous for acquiring global semantic information and robust contextual representations. However, its effectiveness in context compression tasks remains underexplored. Motivated by these findings, we incorporate bidirectional attention into the compressor to enhance contextual modeling during the compression phase.
@@ 3.1 Task Formulation
Additionally, during finetuning, the downstream task loss $\mathcal{L}_{\mathbf{QA}}$ (taking QA as an exemplar downstream task) enhances the ability of $\tilde{M}$ to extract information that is potentially relevant for downstream tasks. For example, for question answering (QA) tasks, the encoder learns through $\mathcal{L}_{\mathbf{QA}}$ to identify and preserve information in the context that is likely to be queried, enabling accurate answer generation $A=(a_{1},a_{2},\ldots,a_{|A|})$ when presented with subsequent questions $Q=(q_{1},q_{2},\ldots,q_{|Q|})$, as shown in Figure 1c; mathematically, $\mathcal{L}_{\mathbf{QA}}=-\log P(A|\tilde{M},Q)$.
@@ 3.2 Semantic-Anchor Compressor
A key distinction between our approach Semantic-Anchor Compression (SAC) and previous methods is that we derive compressed representations directly from selected context tokens, as shown in Figure 2. This involves selecting a subset of tokens $S$ from context $C$ as anchor tokens $S\subseteq C$. We believe that a good selection strategy benefits SAC. Following EPL, our default strategy divides the entire context $C$ into $|S|$ chunks and selects the middle token from each chunk. This setting helps maximize coverage of context $C$. As illustrated in Figure 3a, each selected token $c_{i}\in S$ is enhanced with the anchor embedding $e_{A}$, yielding an embedding sequence $E=(e_{1},e_{2},\ldots,e_{|C|})$:
@@
where $\mathbf{1}_{c_{i}\in S}$ is an indicator function that equals 1 when $c_{i}\in S$ and 0 otherwise. Following previous works, we employ a LLM with LoRA parameters $\theta_{LoRA}$ as the compressor: $\tilde{M}=\mathcal{E}(C)=\mathbf{LLM}(E|\theta_{LoRA})$. Overall, using original tokens from the context avoids learning compression tokens from scratch and we empirically validate that such design improves learning efficiency.
We notice that because the encoder uses causal attention, the anchor tokens $S$ do not have visibility to the full sentence, limiting their representation power. Hence in SAC, we modify the LLM to replace causal attention with bidirectional attention (see Figure 3b), to enhance the LLM’s encoding capability. Notably, this bidirectional attention operates across all tokens, rather than being limited to the anchor tokens, allowing the model to capture richer contextual dependencies.
$\tilde{M}$ can be either the output of anchor tokens from the LLM’s final layer or the key-value pairs from each layer. Following 500xCompressor, we use key-value pairs as the compressed representation $\tilde{M}$. For the encoder, we adopt the same language model as the decoder and leverage the original context’s KV cache, enabling the decoder to comprehend the compressed representation more readily without requiring additional semantic alignment. [^1] For context compression, adopting different LLMs for encoder and decoder will create semantic gaps between the two and will impair performance, as noticed by (Lv et al., 2026). During pretraining, we only use $\mathcal{L}_{\mathbf{LM}}$ and do not use $\mathcal{L}_{\mathbf{AE}}$ to train the compressor. Following previous work, we use $\mathcal{L}_{\mathbf{QA}}$ for finetuning.
@@ 4.1 Experimental Setting
Dataset. For continued pretraining, we use the corpus SlimPajama-6B (Soboleva et al., 2023). During finetuning and evaluation, we employ the standard MRQA (Fisch et al., 2019) question-answering dataset, which consolidates multiple QA tasks and standardizes them into a unified format. We evaluate SAC on both test sets, namely in-domain (ID) and out-of-domain (OOD), to comprehensively assess its in-distribution fitting ability and cross-domain generalization performance.
@@
Implementation Details. SAC utilizes Llama-3.2-1B (Grattafiori and Dubey, 2024) as both the encoder and target LLM. We equip the encoder with trainable LoRA (Hu et al., 2022) adapters (rank = 128, $\alpha=256$), while the target LLM parameters remain frozen. For each context, we partition it into sub-contexts of 510 tokens each. The compressor compresses each sub-context into a sub-compressed representation, and subsequently concatenates these sub-compressed representations to form the complete compressed representation. The number of anchor tokens $|S|=\lfloor L/r\rfloor$ is determined by the compression ratio $r$ and the length of the sub-context $L$. We train all models in two stages: pretraining for 20,000 optimization steps followed by finetuning for an additional 20,000 steps, both conducted with a batch size of 16. Complete hyperparameter configurations are provided in Appendix A.